Let f(x) = 4x^2 – 6 and g(x) = x – 2. Find ( f O g ) (−1).
Answer by Alan3354(69443)

Let f(x) = 4x^2 – 6 and g(x) = x – 2. Find ( f O g ) (-1).
g(-1) = -1 - 2 = -3
f(-3) = 4(-3)^2 - 6 = 36-6
----
f(g(-1)) = 30